Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A computer program generates a single digit by a random process, according to which the probability of generating any

Expert replies
by M7MBA » Sun Jun 14, 2020 1:36 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A computer program generates a single digit by a random process, according to which the probability of generating any digit is directly proportional to the reciprocal of one more than that digit. If all digits are possible to generate, then the probability of generating an odd prime digit is between

A. 0 and 1/6
B. 1/6 and 1/3
C. 1/3 and 1/2
D. 1/2 and 2/3
E. 2/3 and 5/6

[spoiler]OA=B[/spoiler]

Source: Manhattan GMAT
Join the discussion
Source: — Problem Solving |

M7MBA wrote: ↑
Sun Jun 14, 2020 1:36 pm
A computer program generates a single digit by a random process, according to which the probability of generating any digit is directly proportional to the reciprocal of one more than that digit. If all digits are possible to generate, then the probability of generating an odd prime digit is between

A. 0 and 1/6
B. 1/6 and 1/3
C. 1/3 and 1/2
D. 1/2 and 2/3
E. 2/3 and 5/6

[spoiler]OA=B[/spoiler]

Source: Manhattan GMAT
Let's first forget about the condition given in the question and treat this question as

If all digits are possible to generate, then the probability of generating an odd prime digit is between

We know that there are 0-9: ten digits. Out of these 10 digits, there are 3 odd prime digits (3, 5 and 7).

So, the probability of choosing three digits out of 10 digits = 3/10 = 0.3

Looking ar the options, we see that 0.3 lies between 1/6 and 1/3 ( 0.167 and 0.33), so option B seems to be the correct answer.

But is it the correct answer? We must analyze further since 0.3, a probable value is too close to extreme value 0.33 of option B. In the light of option C (0.33 and 0.5), whose extreme value is also 0.33, is a close contender.

Now let's understand the question.

The question means that the probabilities of

1. The digit 0 \( \propto \frac{1}{(0 + 1)}= 1\)
2. The digit 1 \( \propto \frac{1}{(1 + 1)}= \frac{1}{(2)}\)
.
4. The digit 3 \( \propto \frac{1}{(3 + 1)}= \frac{1}{(4)}\)
.
6. The digit 5 \( \propto \frac{1}{(5 + 1)}= \frac{1}{(6)}\)
.
8. The digit 7 \( \propto \frac{1}{(7 + 1)}= \frac{1}{(8)}\)
.
10. The digit 9 \( \propto \frac{1}{(9 + 1)}= \frac{1}{(10)}\)

Note that the above fractions are not the values of probabilities, but the factors to which the probabilities are proportional to.

If we take '5' as the mean value of digits 0 to 9, we can take ignore the impact of the addition of '1' to its reciprocal. But the impacts of the addition of '1' on the reciprocal of '3' is more than that of '7'. So, the combined probability of getting 3, 5, or 7 would be slightly less than 0.3, so the correct answer would be option B.

Correct answer: B

Hope this helps!

-Jay
_________________
Manhattan Review GMAT Prep

Locations: SAT Classes | SAT Tutoring | ACT Tutoring | TOEFL Tutoring | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

M7MBA wrote: ↑
Sun Jun 14, 2020 1:36 pm
A computer program generates a single digit by a random process, according to which the probability of generating any digit is directly proportional to the reciprocal of one more than that digit. If all digits are possible to generate, then the probability of generating an odd prime digit is between

A. 0 and 1/6
B. 1/6 and 1/3
C. 1/3 and 1/2
D. 1/2 and 2/3
E. 2/3 and 5/6

[spoiler]OA=B[/spoiler]

Solution:

Since the sum of the probabilities of generating each single digit is 1, we can create the equation:

k + (1/2)k + (1/3)k + … + (1/10)k = 1

where k, (1/2)k, (1/3)k, …, (1/10)k are the probabilities of generating 0, 1, 2, …, 9, respectively.

We can factor out k from the equation above:

k[1 + 1/2 + 1/3 + … + 1/10] = 1

Let S = 1 + 1/2 + 1/3 + … + 1/10, so we have:

k * S = 1

k = 1/S

Now, let’s find a lower estimate and an upper estimate for S (note: the italic terms are less than the corresponding terms in S, the bold terms are greater and the regular ones are equal):

1/4 + 1/4 + 1/4 + 1/4 + 1/8 + 1/8 + 1/8 + 1/8 + 1/16 + 1/16 < S < 1 + 1/2 + 1/2 + 1/4 + 1/4 + 1/4 + 1/4 + 1/8 + 1/8 + 1/8

1 + 1/2 + 1/8 < S < 1 + 1 + 1 + 3/8

(8 + 4 + 1)/8 < S < (8 + 8 + 8 + 3)/8

13/8 < S < 27/8

Since k = 1/S, we have:

8/27 < k < 8/13

The probability of generating an odd prime digit (i.e., 3, 5, and 7) is k[1/4 + 1/6 + 1/8] = k(13/24). If k is 8/27, then this probability is 8/27 * 13/24 ≈ 1/2 *1/3 = 1/6. If k is 8/19, then this probability is
8/13 * 13/24 = 1/1 * 1/3 = 1/3. We see that the desired probability is between 1/6 and 1/3.

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion